# EarthMarker: Visual Prompt Learning for Region-level and Point-level Remote Sensing Imagery Comprehension

### v2

@@ Abstract
Recent advances in visual prompting in the natural image area have allowed users to interact with artificial intelligence (AI) tools through various visual marks such as box, point, and free-form shapes. However, due to the significant difference between the natural and remote sensing (RS) images, existing visual prompting models face challenges in RS scenarios. Moreover, RS MLLMs mainly focus on interpreting image-level RS data and only support interaction with language instruction, restricting flexibility applications in the real world. To address those limitations, the first visual prompting model named EarthMarker is proposed, which excels in image-level, region-level, and point-level RS imagery interpretation. Specifically, the visual prompts alongside images and text instruction input into the large language model (LLM), adapt models toward specific predictions and tasks. Subsequently, a sharing visual encoding method is introduced to refine multi-scale image features and visual prompt information uniformly. Furthermore, to endow the EarthMarker with versatile multi-granularity visual perception abilities, the cross-domain phased learning strategy is developed, and the disjoint parameters are optimized in a lightweight manner by leveraging both the natural and RS domain-specific knowledge. In addition, to tackle the lack of RS visual prompting data, a dataset named RSVP featuring multi-modal fine-grained visual prompting instruction is constructed. Extensive experiments are conducted to demonstrate the proposed EarthMarker’s competitive performance, representing a significant advance in multi-granularity RS imagery interpretation under the visual prompting learning framework. Our code and dataset are available at https://github.com/wivizhang/EarthMarker.
@@
Visual prompting refers to the technique of guiding the visual models to focus on the region of interest and improving their finer-grained interaction performance by providing them with visual marks (e.g., boxes, points, masks) or examples[1, 2, 3]. Recently, multi-modal large language models (MLLMs) [4, 5] have experienced remarkable advancements in the remote sensing (RS) domain. However, those MLLMs only support language instruction and fail to understand the images in a visual prompting manner. Considering that the high-resolution RS imagery is characterized by scale variation, across categories and tiny objects, fine-grained reasoning is necessary alongside holistic scene interpretation. This is crucial to perform more detailed analyses to make informed decisions in real-world applications[6]. Nevertheless, most existing MLLMs achieve visual-language alignment using image-text pairs, lacking fine-grained referring understanding abilities, such as region-level and point-level. At present, leveraging the visual prompting method to enhance the complex visual reasoning capabilities of MLLMs in RS remains under-explored.
@@
Notably, prompting engineering [7, 8] has been extensively studied in the natural language processing (NLP) community[9] and subsequently spread to the computer vision area. A key example is the Segment Anything (SAM) [10] model, which utilizes multiple visual prompting marks to realize zero-shot segmentation adapted for various new image distributions. However, SAM lacks semantic information, hindering real-world applications. GPT4RoI [11] and RegionBlip [12] have enabled MLLMs to complete region-level visual understanding tasks by training on region-text pairs. Nevertheless, they only support bounding boxes as visual prompts, which lacks flexibility. Osprey [13] excels in pixel-level visual understanding but relies on pre-attached segmentation models, constraining its application range. Additionally, Ferret [14] and SPHINX-V [15] support free-shape visual prompting marks to achieve pixel-level image comprehension. However, all these models are trained on natural scene data, leading to inferior performance when handling the RS imagery.
In the RS field, there are limited works devoted to region-level imagery interpretation. For example, RSVG [16] adopts language prompting to inquire and localize the specific object, but does not involve the visual prompting technique to realize various visual tasks. In addition, inspired by SAM, RSPrompter [17] introduces an automated prompts generation to develop interactive segmentation in RS imagery. Another representative work is EarthGPT [5], which shows the potential of region-level image comprehension by training on visual grounding datasets. However, EarthGPT only supports language interaction without visual prompts, thereby lacking flexibility. These limitations hinder the development of fine-grained spatial understanding and complex reasoning execution in the RS domain. It is clear that the visual prompts learning in the RS domain is still in its infancy.
@@
To bridge this gap, a fine-grained MLLM named EarthMarker is proposed, leveraging visual prompting to extend the capability of MLLMs for region-level and point-level understanding in the RS domain for the first time. Based on the visual prompts learning, as illustrated in Fig. 1, EarthMarker excels at the multi-granularity interpretation of RS imagery across image, region, and point levels. Moreover, EarthMarker can complete a wide range of RS visual reasoning tasks, including scene classification, referring object classification, captioning, inter-relationship analyses, etc. Concretely, the visual prompts, i.e., bounding boxes and points, along with the RS images and the text instructions are provided as input to the LLM. Notably, the visual prompting marks are utilized to isolate specific areas and guide the model to interpret regional content in the entire RS image. Considering that the RS imagery is gathered from an overhead perspective by satellites, associated with large-scale variations and cluttered backgrounds, multi-resolution image input processing is necessary. Subsequently, unlike most existing nature scene visual prompting works, which routinely set different visual encoders and visual prompts encoders. In our method, a sharing visual encoding method is developed. Specifically, the visual prompt is processed to RGB images analogously, which shares the same visual encoder with the inputted image. This strategy is beneficial for consistent feature extraction and understanding the relationship between visual prompts regions and the holistic image, enhancing the performance of the model under visual prompts learning.
@@
In order to enhance the visual prompts-image-text alignment and to equip the EarthMarker with versatile multi-granularity visual comprehension abilities, the cross-domain phased learning strategy is proposed. In the first stage for multi-domain image-text alignment, EarthMarker is trained on the existing nature scene and RS caption data to obtain general image understanding and enhance the modeling of conceptual diversity. Subsequently, the model is further trained on the nature domain referring data to achieve spatial perception in images, beneficial for subsequent developing referring comprehension ability in the RS domain. Lastly, in the RS visual prompting tuning stage, leveraging RS region-text and point-text instruction data, the proposed EarthMarker is equipped with point-level and region-level RS imagery interpretation capability. Notably, the phased training leverages the natural domain generalized knowledge and the RS domain expert knowledge for developing RS visual prompting MLLM. The multi-domain joint training is advantageous for enhancing the deep interpretation of fine-grained RS imagery and improving open-vocabulary reasoning capabilities. In addition, the updatable parameters of the model are disjoint, preventing interference between understanding images at different granularity and the capability to follow the visual prompt instruction.
@@
Another challenge lies in the datasets, e.g., existing visual prompting datasets [14, 13] are restricted to the natural scene, lacking RS semantics. It has become indispensable to construct a visual prompting dataset tailored to the RS domain for developing fine-grained MLLM. To this end, a RS visual prompting dataset named RSVP-3M, featuring large-scale fine-grained instruction-following, is developed. In particular, diverse publicly available RS data are transformed and re-annotated into uniform conversation formats. Furthermore, part of the more high-quality caption data is generated from GPT-4V[18]. Those captions are uniquely tailored with the distinctive characteristics of each RS imagery, thereby enhancing the richness and diversity of data. Through the data conversion and re-annotation from existing datasets and GPT-4V, over 3M image-point-text and image-region-text pairings are constructed, covering a wide geographic distribution and multiple types of ground targets.
@@
The First RS Visual Prompting Dataset, RSVP. A large-scale RS regional instruction dataset named RSVP-3M, containing over 3M image-point-text and image-region-text pairings, is constructed. The construction of RSVP-3M facilitates fine-grained RS imagery interpretation, laying the foundation for the development of visual prompting in the RS domain.
The First RS Visual Prompting MLLM, EarthMarker. Leveraging our newly constructed RSVP, the visual prompting MLLM named EarthMarker is proposed. EarthMarker can interpret RS imagery in the multi-turn conversation at different granularity, including image, region, and point levels, significantly catering to the fine-grained interpretation needs for RS imagery.
The First RS Visual Prompt Learning Framework. A universal region and point-level visual prompting data annotation method is developed. Subsequently, a sharing visual encoding mechanism is proposed, which adapts visual prompts to match the dimensions of the input image, thereby both of them undergo uniform processing by the same visual encoder. This mechanism comprehensively enhances the interplay among visual prompts, holistic images, and text instructions. Furthermore, the cross-domain phased learning strategy is designed, and the disjoint parameters are optimized in a lightweight manner by leveraging the multi-domain data, endowing EarthMarker with spatial perception and visual prompting following capabilities.
Superior Performance on Multi-granularity RS Visual Tasks. Extensive experiments are conducted to demonstrate EarthMarker’s competitive performance in multi-granularity RS visual interpretation tasks, compared with the SOTA specialist models, MLLMs, and visual prompting models. The tasks evaluated include scene classification, referring object classification, captioning, and inter-relationship analyses. Therefore, EarthMarker successfully explores the adaptation of the visual prompt learning in the RS domain, improving the performance of MLLM and representing a significant step in fine-grained RS imagery interpretation.
@@ II-A Multi-modal Large Language Models (MLLMs)
Recently, the advancement of large language models (LLMs) has significantly fueled the revolution and innovation in the natural language processing (NLP) field. The representative works including closed-source GPT series [20, 11] and open-source LLaMA series [21, 22] have achieved powerful generalizable language processing and reasoning ability. Inspired by LLM and by further injecting visual signals, MLLMs are developed for visual-language mutual comprehension and various visual tasks. For example, VisualGPT [23], BLIP [24] and Flamingo [25] show strong multi-modal reasoning potential after aligning LLMs with visual modality. Notably, LLAMA-Adapter V2 [26] and SPHINX [27] adopt zero-shot attention mechanism and linear projection layers tuning to mix LLM with visual signal. Those nature scene MLLMs laid the foundation for the extension to the remote sensing (RS) domain.
Some pioneer RS MLLMs have emerged, and related studies such as EarthGPT [5], Geochat [4], and SkyEyeGPT [28] have enabled MLLMs to interpret RS imagery. Among them, Geochat is the first MLLM targeting solving multiple tasks on optical RS images. Furthermore, EarthGPT has proposed a more universal MLLM that can deal with multi-source RS imagery and a wide range of RS visual tasks. There is no doubt that those models facilitate the development of MLLMs in the RS-specific domain. However, those models complete visual interpretation only through human-like language interactions, but cannot generate responses through visual prompts. Apparently, existing RS MLLMs mainly focus on image-level and visual grounding, but are incapable of referring comprehension. Therefore, this paper aims to enhance the MLLMs for referring fine-grained understanding of vision.
@@ II-B Prompt Engineering
Prompt engineering is an emerging research direction in NLP [20]. Representation works contain AutoPrompt [7] and CoOp [8], which are designed to automate prompt template generation for language and vision-language models, instead of manual crafting. Additionally, language prompting has been applied for developing open-vocabulary detection models such as DetPro [29] and Promptdet [30]. Compared with the extensively developed language prompting technique, visual prompting also needs more exploration. A major development is the Segment Anything (SAM) [10] model, which supports multiple segmentation prompts to enhance the zero-shot performance. Due to the lack of semantic labels in SAM, the Semantic-SAM [31] is proposed to realize multi-level semantics analysis and prediction. Notably, GPT4RoI [11] uses spatial boxes, and combines language and region-of-interest for input, enabling regional recognition. Colorful Prompting Tuning (CPT) [32] uses color-based markers to improve the performance of pre-trained vision-language models. The aforementioned models are trained on nature scene datasets. Note that Osprey [13] incorporates fine-grained mask regions into language instruction, achieving pixel-level visual understanding. Other visual prompting works including RegionBlip [12], Kosmos-2 [33], Shikra [34], and Ferret [14], also have shown promising results in region-based image understanding by leveraging visual prompting techniques. Additionally, the study entitled “Visual Prompting via Image Inpainting”[1] shows that various vision tasks can be accomplished well by giving desired task examples.
There are pioneering studies in the RS domain on region-level image understanding. For example, RSVG [16] can provide the referred object’s bounding box based on images and natural language expression. Moreover, EarthGPT [5] also has the visual grounding ability, and it is capable of providing captions for specific areas within images. Inspired by prompt learning, RSPrompter [17] designs an automated approach to generate appropriate prompts for SAM input, facilitating RS imagery segmentation. However, RSVG adopts language prompting but without visual prompting, whilst RSPrompter is only tailored to the segmentation task. Apparently, there is no unified visual prompting framework designed for the RS domain to further improve the performance of MLLMs. Those limitations hamper the development of more complex and fine-grained RS imagery understanding, therefore this paper focuses on filling this gap.

### v3

# EarthMarker: A Visual Prompting Multi-modal Large Language Model for Remote Sensing

@@ Abstract
Recent advances in prompt learning have allowed users to interact with artificial intelligence (AI) tools in multi-turn dialogue, enabling an interactive understanding of images. However, it is difficult and inefficient to deliver information in complicated remote sensing (RS) scenarios using plain language instructions alone, which would severely hinder deep comprehension of the latent content in imagery. Besides, existing prompting strategies in natural scenes are hard to apply to interpret the RS data due to significant domain differences. To address these challenges, the first visual prompting-based multi-modal large language model (MLLM) named EarthMarker is proposed in the RS domain. EarthMarker is capable of interpreting RS imagery at the image, region, and point levels by levering visual prompts (i.e., boxes and points). Specifically, a shared visual encoding method is developed to establish the spatial pattern interpretation relationships between the multi-scale representations of input images and various visual prompts. Subsequently, the mixed visual-spatial representations are associated with language instructions to construct joint prompts, enabling the interpretation of intricate content of RS imagery. Furthermore, to bridge the domain gap between natural and RS data, and effectively transfer domain-level knowledge from natural scenes to the RS domain, a cross-domain learning strategy is developed to facilitate the RS imagery understanding. In addition, to tackle the lack of RS visual prompting data, a dataset named RSVP featuring multi-modal multi-granularity visual prompts instruction-following is constructed. Extensive experiments are conducted to demonstrate the competitive performance of the EarthMarker. The proposed EarthMarker represents a significant advance in multi-granularity RS imagery interpretation under the visual prompting learning framework. Our code and dataset are available at https://github.com/wivizhang/EarthMarker.
@@
Visual prompting is a current research hotspot in natural domains, aiming to interpret regions of interest in images with the guidance of visual marks (e.g., boxes, points, and scribbles)[1, 2]. Particularly, injecting visual prompts into multi-modal large language models (MLLMs) can establish a joint prompting mechanism, which allows leveraging both visual and textual instructions. The joint prompting is a human-like expression way, enabling a more flexible and efficient interactive manner to understand images. Applying visual prompting to the remote sensing (RS) domain must account for the unique characteristics of RS imagery, which features scale variation, cross-category diversity, and complex contextual semantic information [3, 4]. Notably, a major challenge lies in effectively guiding models to recognize and interpret critical latent information from complicated RS imagery. However, it is difficult to efficiently describe the specific regions in complex RS images using plain language alone [5, 6]. Besides, existing MLLMs mainly achieve image-level visual-language alignment, which results in the restricted ability for detailed image understanding[7]. Thus, it is indispensable to develop more concise and effective prompting mechanisms to enhance fine-grained visual reasoning in the RS domain.
@@
Notably, prompting engineering [8, 9] has been extensively studied in the natural language processing (NLP) community[10] and subsequently spread to the computer vision area. A key example is the Segment Anything (SAM) [11] model, which utilizes multiple visual prompts to realize zero-shot segmentation adapted for various new image distributions. GPT4RoI [12] and RegionBlip [13] have enabled MLLMs to complete region-level visual understanding tasks by training on region-text pairs. Osprey [14] excels in pixel-level visual understanding but relies on pre-attached segmentation models, constraining its application range. Additionally, Ferret [15] and SPHINX-V [16] support free-shape visual prompting marks to achieve pixel-level image comprehension. Nonetheless, all these models are trained on natural scene data, leading to inferior performance when handling the RS imagery.
In the RS field, there are limited works devoted to region-level fine-grained image understanding or visual prompting. For example, RSVG [17] adopts language prompting to inquire and localize the specific object. Other representative RS MLLMs include EarthGPT [18] and Geochat [6], which extend beyond image-level understanding to achieve region-level visual perception through training on visual grounding data. In addition, SAM also shows limited effectiveness on RS imagery since its ViT backbone is pre-trained on large-scale close-range RS datasets[19, 20]. Inspired by SAM, RSPrompter [21] introduces an automated prompts generation to develop interactive segmentation tailored for RS data. However, these RS MLLMs primarily realize image-text alignment and rely solely on language instructions, which struggle to uncover valuable information hidden within complex RS scenes and lack interactive flexibility. Furthermore, current visual prompting models including SAM and its variants are limited to image segmentation, lacking the ability for visual reasoning of complicated imagery. Therefore, the integration of visual prompting into MLLMs for detailed interpretation of RS data remains under-explored.
@@
To bridge the gap, a visual prompting MLLM named EarthMarker is proposed, leveraging visual prompts to extend the capability of MLLMs for region-level and point-level understanding in the RS domain for the first time. As illustrated in Fig. 1, EarthMarker excels at the multi-granularity interpretation of RS imagery. Specifically, EarthMarker can switch levels to interpret an image, ranging from coarse-grained holistic scene-level to fine-grained region/object-level, and even point-level distinctions. Moreover, EarthMarker can complete a wide range of both coarse-grained and fine-grained RS visual tasks, including scene classification, referring object classification, image/region/point-level captioning, inter-relationship analyses, etc. Notably, the visual prompts are used to isolate specific areas and guide the model to interpret regional content in the entire RS image. Concretely, the visual prompts, i.e., bounding boxes and points, along with the RS images and the text instructions are provided as input to the large language model (LLM), adapt models toward specific predictions and tasks.
In fact, there is a considerable gap between the different modalities involved in sparse visual prompts, dense RS image features, and textual features that operate on a semantic level. The significant differences in feature richness and structure make it challenging to align and integrate these modalities. To address this problem, in our method, a sharing visual encoding method is developed. Specifically, the visual prompt is processed to RGB images analogously, which shares the same visual encoder with the input image. This strategy is beneficial for consistent feature extraction and understanding the relationship between visual prompts regions and the holistic image, enhancing the performance of the model under visual prompts learning. The designed joint prompting mechanism, combining visual prompts and text instructions, enables EarthMarker to dynamically adapt to diverse tasks, interpretation granularity, and varying spatial resolution imagery.
@@
The limited amount of RS data hinders achieving region and text mutual interaction, understanding spatial relationships, and object localization. To mitigate this limitation, general-domain data is utilized for mixed training. Most importantly, to enhance the visual prompts-image-text alignment, the cross-domain learning strategy is proposed. Specifically, in the first stage for multi-domain image-text alignment, EarthMarker is trained on the existing nature scene and RS caption data to obtain general image understanding and enhance the modeling of conceptual diversity. Subsequently, the model is further trained on the nature domain referring data to achieve spatial perception in images, beneficial for subsequent developing referring comprehension ability in the RS domain. Lastly, in the RS visual prompting tuning stage, we leverage RS region-text and point-text instruction data, which features a variety of spatial resolutions, for domain adaption training. The proposed EarthMarker is equipped with holistic and point/region-level RS imagery interpretation capabilities. Notably, the cross-domain learning training leverages the natural domain generalized knowledge and the RS domain expert knowledge for developing RS visual prompting MLLM. The multi-domain joint training is advantageous for enhancing the deep interpretation of fine-grained RS imagery and improving open-vocabulary reasoning capabilities. In addition, the updatable parameters of the model are disjoint, preventing interference between understanding images at different granularity and the capability to follow the visual prompt instruction.
@@
Another challenge lies in existing visual prompting datasets [15, 14] are restricted to the natural scene, lacking RS semantics. Moreover, current public RS captions data tend to be overly simplistic and repetitive, without unique characteristics from the RS imagery. It has become indispensable to construct a visual prompting dataset tailored to the RS domain. To this end, an RS visual prompting dataset named RSVP, featuring multi-modal large-scale visual-language joint instruction-following, is developed. In particular, diverse publicly available RS data are transformed and re-annotated into uniform conversation formats. Furthermore, part of the more high-quality caption data is generated from GPT4V[22]. Those captions are uniquely tailored with the distinctive characteristics of each RS imagery, thereby enhancing the richness and diversity of data. Through the data conversion and re-annotation from existing datasets and GPT4V, approximately 3.65 M image-point-text and image-region-text pairings are constructed. In addition, RSVP includes images with a variety of spatial resolutions, covering a wide geographic distribution. These characteristics of the RSVP dataset equip EarthMarker to effectively handle both fine details and broader patterns, enabling it to generalize well in real-world applications.
@@
First RS Visual Prompting MLLM, EarthMarker. A visual prompting MLLM called EarthMarker is proposed in the RS domain for the first time. EarthMarker can comprehend RS imagery under visual and text joint prompts, and flexibly switch interpretation levels, including image, region, and point levels. More importantly, the proposed EarthMarker fills the gap in visual prompting MLLMs for RS, significantly catering to the fine-grained interpretation needs of RS imagery in real-world applications.
First RS Visual Prompt Learning Framework. A universal region-level and point-level visual prompting data annotation method is developed. Subsequently, a sharing visual encoding mechanism is proposed to enhance the interplay among visual prompts, holistic images, and text instructions. Furthermore, the cross-domain learning strategy is designed, and the disjoint parameters are optimized in a lightweight manner by leveraging the multi-domain data, endowing EarthMarker with spatial perception and joint instruction-following capabilities.
First RS Visual Prompting Dataset, RSVP. A large-scale RS regional instruction dataset named RSVP, containing roughly 3.65 M image-point-text and image-region-text pairings, is constructed for the first time. The construction of RSVP facilitates fine-grained RS imagery interpretation, laying the foundation for the development of visual prompting in the RS domain.
Superior Performance on Multi-granularity RS Visual Tasks. Extensive experiments are conducted to demonstrate EarthMarker’s competitive performance in multi-granularity RS visual interpretation tasks, compared with the SOTA specialist models, MLLMs, and visual prompting models. The tasks evaluated include scene classification, referring object classification, region captioning, inter-relationship analyses, etc. Therefore, EarthMarker successfully explores the adaptation of the visual prompt learning in the RS domain, improving the performance of MLLM and representing a significant step in fine-grained RS imagery interpretation.
@@ II-A Multi-modal Large Language Models (MLLMs)
Recently, the advancement of large language models (LLMs) has significantly fueled the revolution and innovation in the natural language processing (NLP) field. The representative works including closed-source GPT series [24, 12] and open-source LLaMA series [25, 26] have achieved powerful generalizable language processing and reasoning ability. Inspired by LLM and by further injecting visual signals, MLLMs are developed for visual-language mutual comprehension and various visual tasks. For example, VisualGPT [27], BLIP [28] and Flamingo [29] show strong multi-modal reasoning potential after aligning LLMs with visual modality. Notably, LLAMA-Adapter V2 [30] and SPHINX [31] adopt zero-shot attention mechanism and linear projection layers tuning to mix LLM with visual signal. Additionally, VisionLLM[32] aligns the definitions of vision-centric tasks with LLM methodologies to address vision-centric tasks in an open-ended manner. Those nature scene MLLMs laid the foundation for the extension to the RS domain.
Some pioneer RS MLLMs have emerged, and related studies such as EarthGPT [18], Geochat [6], SkyEyeGPT [33], and LHRS-Bot[34] have enabled MLLMs to interpret RS imagery. Among them, Geochat is the first MLLM targeting solving multiple tasks on optical RS images. EarthGPT has proposed a universal MLLM that can deal with multi-source RS imagery and a wide range of RS visual tasks. LHRS-Bot constructs the LHRS-Bench, to facilitate the RS community in evaluating the RS-specific MLLMs across diverse evaluation dimensions. There is no doubt that those models facilitate the development of MLLMs in the RS-specific domain. However, those models complete visual interpretation only through language interactions, but cannot generate responses guided by visual prompts to handle referring visual comprehension tasks. Apparently, existing RS MLLMs mainly focus on image-level and visual grounding, but are incapable of more fine-grained interpretation across region, point, and pixel levels. Therefore, this paper aims to enhance the MLLMs for the versatile multi-granularity understanding of RS imagery.
@@ II-B Prompt Engineering
Prompt engineering is an emerging research direction in NLP [24]. Representation works contain AutoPrompt [8] and CoOp [9], which are designed to automate prompt template generation for language and vision-language models, instead of manual crafting. Additionally, language prompting has been applied for developing open-vocabulary detection models such as DetPro [35] and Promptdet [36]. Compared with the extensively developed language prompting technique, visual prompting also needs more exploration.
A major development is the SAM [11] model, which supports multiple segmentation prompts to enhance the zero-shot performance. Due to the lack of semantic labels in SAM, the Semantic-SAM [37] is proposed to realize multi-level semantics analysis and prediction. Notably, GPT4RoI [12] uses spatial boxes, and combines language and region-of-interest for input, enabling regional recognition. Colorful Prompting Tuning (CPT) [38] uses color-based markers to improve the performance of pre-trained vision-language models. Note that Osprey [14] incorporates masked regions into language instruction, achieving pixel-level visual understanding. Other visual prompting works including RegionBlip [13], Kosmos-2 [39], Shikra [40], and Ferret [15], also have shown promising results in region-based image understanding by leveraging visual prompting techniques. The study entitled “Visual Prompting via Image Inpainting”[41] shows that various vision tasks can be accomplished well by giving desired task examples. Furthermore, RegionPLC[42] has proposed a lightweight and scalable regional point-language contrastive learning framework for open-world 3D scene understanding. However, the aforementioned models are trained on natural datasets, and the significant differences between natural and RS imagery suggest that their direct applicability to RS tasks is limited.
Limited studies in the RS domain have emerged on language or visual instruction learning recently, besides traditional vision models[43, 44]. For example, RSVG [17] can provide the referred object’s locations based on natural language expression. EarthVQA[45] dataset alongside a semantic object awareness framework (SOBA) is developed. EarthVQA contains images, corresponding semantic visual prompts, and QA pairs for training SOBA to achieve object-centric reasoning. Inspired by prompt learning, RSPrompter [21] designs an automated approach to generate appropriate prompts for SAM input, facilitating RS imagery segmentation. However, RSVG adopts language prompting but without visual prompting. EarthVQA solely involves pixel-level reasoning on city planning images, and RSPrompter is only tailored to prompt-based segmentation tasks. Additionally, there are expert models that have developed detection and semantic segmentation with point supervision, including PSOD[46, 47, 48, 49]. Nevertheless, those models are designed for single tasks without versatile capabilities. Obviously, there is no unified visual prompting framework designed for the RS domain to further improve the multi-granularity interpretation abilities of MLLMs. Those limitations hamper the development of more complex and fine-grained RS imagery understanding, therefore this paper focuses on filling this gap.